Parth (parthkohli):

8 men and 6 women do some work in 30 days; 14 men and 10 women do the same work in 20. How many days will it take 5 men and 5 women? (assume that men and women take unequal amounts of time)

Parth (parthkohli):

Ugh, Dad gives hard questions.

Parth (parthkohli):

Though I know that this is related to inverse relations and all. I know how to do questions for one guy, but not for so many.

OpenStudy (asnaseer):

I can guide you on how to do this when you are online.

OpenStudy (phi):

Think in terms of rate * time = job rate is measured in \[\frac{\text{jobs}}{\text{man-days}} \] set up 2 equations and 2 unknown rates (for men and women). For example, the first equation is \[ (8 \text{ men} \cdot m \frac{\text{jobs}}{\text{man-days}} + 6 \text{ women} \cdot w \frac{\text{jobs}}{\text{woman-days}})\cdot 30 \text{ days}= 1 \text{ job} \] Once you find the rates m and w, solve for D in \[ (5m+5w)*D \text{ days} = 1 \text{ job} \] The peculiar thing about this problem is that one of the rates is negative! (slows down the work)

Parth (parthkohli):

@asnaseer Hello!

OpenStudy (asnaseer):

hi - looks like phi has put down some explanations already - do you understand them?

OpenStudy (asnaseer):

I tend to do these in a /slightly/ different manner

Parth (parthkohli):

Okay, please, if you can. :)

OpenStudy (asnaseer):

ok, you are told that "8 men and 6 women do some work in 30 days" this means that the work can be done in 1 day if we have "30(8 men and 6 women)" agreed?

OpenStudy (asnaseer):

30 times the work force means it is completed in 1/30'th of the time

Parth (parthkohli):

Yes, agreed. Sorry... there are some problems in the internet at my end. Please don't mind that. Thank you :)

OpenStudy (asnaseer):

ok, similarly you are told that "14 men and 10 women do the same work in 20" this means that the work can be done in 1 day if we have "20(14 men and 10 women)" agreed?

Parth (parthkohli):

Yes, agreed again!

OpenStudy (asnaseer):

ok, next I will represent the "work efficiency" of a man by "m" and that of a women by "w". this gives us: 30(8m + 6w) = 20(14m + 10w)

Parth (parthkohli):

OK, I am following. Yes.

OpenStudy (asnaseer):

we can divide both sides by 10 to get: 3(8m + 6w) = 2(14m + 10w)

OpenStudy (asnaseer):

which leads to: w = -2m

Parth (parthkohli):

Yes... and substitution now, right?

Parth (parthkohli):

\[3(8m + 6(-2m)) = 2(14m + 10(-2m)) \]Yes?

OpenStudy (asnaseer):

yes, now substitute this into one of the equations, say this one: 30(8m + 6w) = 30(8m - 12m) = 30 * (-4m) = -120m

OpenStudy (asnaseer):

that represents the total work units

OpenStudy (asnaseer):

so 5 men and 5 women would have total work units of: 5(m + w) = 5(m - 2m) = -5m

OpenStudy (asnaseer):

therefore total days taken will be: days = (-120m)/(-5m)

OpenStudy (asnaseer):

hope that makes sense?

OpenStudy (asnaseer):

if not, try using the method that phi described. use the one that makes most sense to you.

Parth (parthkohli):

Yes!

OpenStudy (asnaseer):

gr8! :)

Parth (parthkohli):

Thank you so much, @asnaseer!

OpenStudy (asnaseer):

yw :)